8. Problems 2.8 Suppose that a firm has a marginal cost function given by MC(ą)1. Which of the following represent the firms total cost function? (Hint: K is a constant of integration) Fixed costs are represented by As you may know from an earlier economics course, if a firm takes price (p) as given in its decisions then it will produce that output for which p = MC(q). Suppose the price is 15 p = 15 If the firm follows this profit maximizing rule, it will produce q*- fixed cost of K . If the firm is breaking even, with 0 profit, then the firm is incurring a Assume the price p increases to 20, while fixed costs remain the same After the price increase, profits will increase to Assuming the firm sets p-MC(q) to maximize profits, which of the following formulas represents profits as a function of p only? p 98 (p-1)2 - 98 Bl-98 р-20 Integrating the inverse marginal cost function MC-T (p) = p-1 from p = 15 top = 20 yields p= 15
Answer #1

2 o2 Fixed coots ane epreaented 64 K 15 - Ly 儿 K. - 310.98

4 19 2. &OX 19-180.5-19-98 82.5 2 2 2

MC is the derivative of TC.

Profit is the difference between TR and TC.

TR = Pq

At equilibrium; P = MC for this we get value of equilibrium quantity.

Integration of p - 1 = p2/2 - p = (20x20)/2 - 20 = 200 - 20 = 180